the task is 3/7 = 2/x-4
i solved as follows: 0=(2/x-4)/(3/7) = 2/(x-4)*(7/3) now the result would be 14/3x-12 but I know that this is the wrong answer... where is my mistake?

Answer by Alan3354(69443)

the task is 3/7 = 2/x-4
i solved as follows: 0=(2/x-4)/(3/7) = 2/(x-4)*(7/3) now the result would be 14/3x-12 but I know that this is the wrong answer... where is my mistake?
---------------
You divided by (3/7), which makes it = 1, not zero.
3/7 = 2/x-4
Cross multiply
3*(x-4) = 14
3x-12 = 14
3x = 26
x = 26/3